2+3=10
7+2=63
6+5=66
8+4=84
9+7=144
I need help finding the equation?

Answer by graphmatics(170)

I don't know where you actually got THIS problem but I know the solution. What you are looking at is a two variable function, a f(x,y) function. When you look at the numbers it is clear what the expression of the function is.
f(x,y) = x*x + x*y
f(7,2) = 7*7 + 7*2 = 49 + 14 = 63